# EarthMarker: Visual Prompt Learning for Region-level and Point-level Remote Sensing Imagery Comprehension

### v2

@@ V-F Complex Visual Reasoning
In this part, we present the qualitative experimental result of EarthMarker to demonstrate its proficiency in completing complex RS tasks such as key target inter-relationship analyses. The text instruction for the relationship analyses task is “Please analyze the relationship between all marked regions in the image.” As shown in Fig. 5, when faced with an airport scenario, four visual models provided different responses.
@@
Specifically, the response generated by GPT-4V does not specify the exact categories of each marked region. Additionally, GPT4V incorrectly describes Region 1 as “Region 1 supports all other regions indirectly by housing vehicles and equipment necessary for ground operations, including baggage handling, maintenance, and possibly emergency services”, whereas Region 1 actually contains fuel storage tanks. The Vip-LLava model incorrectly identifies all regions as airplanes. Note that Sphinx-V correctly identifies the types of objects in each region and analyzes the internal relationship of some regions. For example, the Sphinx-V answer “They are positioned in close proximity to each other, suggesting they are within the same area of the airport. The tanks are similar in appearance and likely serve the same function.” However, it fails to provide a comprehensive analysis of the relationships among all four regions, and incorrectly states that the airplane in Region 4 is in motion.
@@
In contrast, our EarthMarker delivers an exemplary response. It firstly summarizes the relationship of all marked regions representing different elements of an airport environment. It then analyzes functionally similar areas in detail, stating that “ and are both storage tanks, likely used for fuel or other liquids, and are similar in shape and size, suggesting they are part of the same system or facility. and are both commercial airplanes, indicating the area is used for aircraft operations. The positioning of the planes and the tanks suggests a functional airport with operational infrastructure for aircraft and fuel services.” This response accurately reflects the diverse functionalities within the airport, demonstrating superior comprehension and analysis compared to the other visual prompting MLLMs.
VI Conclusion
@@
In this paper, the fine-grained MLLM called EarthMarker, the first visual prompting model specifically designed for the RS domain, is proposed. Furthermore, the RS visual prompting instruction dataset called RSVP is constructed for the first time, facilitating the development of fine-grained RS imagery comprehension. In addition, the visual prompts learning framework is developed. Particularly, the shared visual encoding method is developed to uniformly refine multi-scale visual features and visual prompt content, which is beneficial for comprehensively understanding the interplay between visual prompts and the holistic image. Subsequently, the referring areas in the input are replaced by the proposed hybrid representation before being fed into the LLM to instruct the model to comprehend referring areas and toward specific predictions. Employing the RSVP-3M and the visual prompt learning framework, EarthMarker is equipped with multi-granularity visual understanding capability at the image, region, and point levels, making it able to simultaneously perform comprehensive and intelligent analysis in real-world scenarios. In the future, we plan to incorporate a broader range of visual modalities into EarthMarker, enhancing its multi-source imagery comprehension capabilities. In addition, we plan to support free-form shapes as visual marks to adjust the referring granularity flexibly.

### v3

@@ VI-B2 Relationship Analyses Tasks
In this part, we present the qualitative experimental result of EarthMarker to demonstrate its proficiency in completing complex RS tasks such as key object inter-relationship analyses. The text instruction for the relationship analyses task is “Please analyze the relationship between all marked regions in the image.” As shown in Fig. 6, when faced with an airport scenario, four visual models provided different responses.
@@
Specifically, the response generated by GPT4V does not specify the exact categories of each marked region. Additionally, GPT4V incorrectly describes Region 1 as “Region 1 supports all other regions indirectly by housing vehicles and equipment necessary for ground operations, including baggage handling, maintenance, and possibly emergency services”, whereas Region 1 actually contains fuel storage tanks. The Vip-LLava model incorrectly identifies all regions as airplanes. Note that Sphinx-V correctly identifies the types of objects in each region and analyzes the internal relationship of some regions. For example, the Sphinx-V answer “They are positioned in close proximity to each other, suggesting they are within the same area of the airport. The tanks are similar in appearance and likely serve the same function.” However, it fails to provide a comprehensive analysis of the relationships among all four regions, and incorrectly states that the airplane in Region 4 is in motion.
@@
In contrast, our EarthMarker delivers an exemplary response. It firstly summarizes the relationship of all marked regions representing different elements of an airport environment. It then analyzes functionally similar areas in detail, stating that “ and are both storage tanks, likely used for fuel or other liquids, and are similar in shape and size, suggesting they are part of the same system or facility. and are both commercial airplanes, indicating the area is used for aircraft operations. The positioning of the planes and the tanks suggests a functional airport with operational infrastructure for aircraft and fuel services.” This response accurately reflects the diverse functionalities within the airport, demonstrating superior comprehension and analysis compared to the other visual prompting MLLMs, which plays a potential key role in solving air-to-ground safety monitoring problems[103].
VI-C Failure Cases
The aforementioned visualization examples indicate that our proposed EarthMarker possesses strong multi-granularity (e.g., image, region, and point-level) visual reasoning capabilities. Compared to other MLLMs and visual prompting models, EarthMarker demonstrates greater potential for real-world applications. However, in some complex situations and difficult conditions, there are still challenges in accurately interpreting small objects in specific regions. For example, as shown in Fig. 7, EarthMarker has limitations in the accuracy of referring object classification for small objects. In the first image, which covers a wide geographic area, we mark two regions: Region 1 (Expressway-Toll-Station), and Region 2 (Overpass). EarthMarker correctly identifies Region 2 but fails to accurately recognize Region 1. In contrast, Sphinx [31] successfully interpreters “Overpass”, while EarthGPT provides a close answer, identifying it as a “Highway”. Other models fail to identify the region correctly and exhibit hallucinations. In the second image, all MLLMs [6, 18, 31] and visual prompting models[100, 16], including EarthMarker, struggle to make correct determinations when faced with tiny objects, which occupy only a few pixels in the image. In summary, while the proposed EarthMarker successfully explores the adaptation of visual prompt learning for multi-granularity RS imagery understanding in the RS domain, there is still significant room for development. In the future, we plan to improve EarthMarker’s abilities in visual understanding of tiny objects under challenging scenarios.
VII Conclusion
@@
In this paper, the first visual prompting model called EarthMarker, specifically designed for the RS domain, is proposed. Moreover, the RS visual prompt-based instruction dataset called RSVP is constructed for the first time, containing roughly 3.65 M image-point-text and image-region-text pairings. In addition, the visual prompt learning framework is developed. Particularly, the shared visual encoding method is developed to uniformly refine multi-scale visual features and visual prompt content, which is beneficial for comprehensively understanding the interplay between local visual prompts and the holistic image. Employing the newly construed RSVP and the visual prompt learning framework, EarthMarker is equipped with multi-granularity visual understanding capability across the image, region, and point levels. The proposed EarthMarker advances the development of fine-grained RS imagery comprehension, providing a comprehensive and intelligent analysis in real-world scenarios. To enhance all-purpose capabilities of EarthMarker, in the future, we plan to incorporate a broader range of various modalities into EarthMarker, enhancing its multi-source, varying spectral imagery and temporal data comprehension capabilities. Furthermore, we plan to support free-form shapes as visual prompts to adjust the interpretation granularity flexibly.